# Exploiting Observation Bias to Improve Matrix Completion

### v2

@@ Appendix B Clustering
$\mathcal{E}_{1}$ =$=\left\{\sum_{k=1}^{n}\mathds{1}\{\mathbf{u}_{k}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i})\}\geq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right\},$ (73)
$\mathcal{E}_{2}$ =$=\left\{\sum_{\ell=1}^{n}\mathds{1}\{\mathbf{v}_{\ell}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j})\}\geq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right\},$ (74)
and note that according to Lemma B.1 and union bounding over the two events we have
@@
$\left\lVert\mathbf{u}_{k}-\mathbf{v}_{l}\right\rVert_{2}\leq\left\lVert\mathbf{z}_{i}-\mathbf{z}_{j}\right\rVert_{2}+\left\lVert\mathbf{u}_{k}-\mathbf{z}_{i}\right\rVert_{2}+\left\lVert\mathbf{v}_{l}-\mathbf{z}_{j}\right\rVert_{2}\leq\sqrt{2}+(2+c)\varepsilon_{n}\leq\sqrt{3}.$ (75)
@@
$\mathcal{E}_{3}=\left\{\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}(\mathbf{z}_{i},\mathbf{z}_{j})\right\rvert\geq\frac{n^{2}\rho_{n}}{512}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right\},$ (81)
@@
[…]
@@
$\rho_{n}n^{2}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\geq 2048\left(\log\left(\frac{1}{\delta}\right)+2d\log\left(\frac{3}{\varepsilon_{n}}\right)\right).$ (83)
@@
$\mathbb{P}\left(\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}(\mathbf{z}_{i},\mathbf{z}_{j})\right\rvert<\frac{\rho_{n}n^{2}}{512}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right)\leq\mathbb{P}\left(\mathcal{E}_{3}^{c}|\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)+\mathbb{P}(\mathcal{E}_{1}^{c}\cup\mathcal{E}_{2}^{c})\leq 3\delta.$ (87)
Appendix C Denoising
Here we argue that every revealed entry of the unfilled clustered outcome matrix $\bar{\mathbf{H}}$ is an accurate estimate of its corresponding entry in $\mathbf{H}$.
Lemma C.1 (Denoising).
Suppose we choose
$N_{n}=cn^{2}\rho_{n}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2},$
for some universal constants $c>0$ large enough. Then, for all $\delta\in(0,1)$, the following event: for every revealed entry $(i,j)$ of $\overline{\mathbf{H}}$
$\left\lvert\bar{h}_{ij}-h_{ij}\right\rvert\leq 14LB\varepsilon_{n}+\sqrt{\frac{\sigma^{2}}{2cn^{2}\rho_{n}}\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\log\left(\frac{2n^{2}}{\delta}\right)}$ (88)
holds with probability at least $1-\delta$, provided that
$\varepsilon_{n}\geq\frac{C}{\rho_{n}^{2}}\sqrt{\frac{d^{3}}{n}\log\left(\frac{e\,n}{\delta}\right)},$ (89)
$n\geq Cd\log\left(\frac{e\,n}{\delta}\right),$ (90)
for some universal constant $C>0$.
Proof of Lemma C.1.
We can express the difference between $\bar{h}_{ij}$ and $h_{ij}$ for any $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ as
$\bar{h}_{ij}-h_{ij}$ $=\underbrace{\frac{1}{|\Omega_{ij}|}\sum_{(k,l)\in\Omega_{ij}}\left(f(\mathbf{u}_{k},\mathbf{v}_{l})-f(\mathbf{u}_{i},\mathbf{v}_{j})\right)}_{\Delta_{ij}}+\underbrace{\frac{1}{|\Omega_{ij}|}\sum_{(k,l)\in\Omega_{ij}}\xi_{kl}}_{\nu_{ij}}.$ (91)
(1) Bounding $\nu_{ij}$. Recall that an entry of $\overline{\mathbf{H}}$ is considered revealed if $\lvert\Omega_{ij}\rvert\geq N_{n}=c\,n^{2}\rho_{n}\left(\varepsilon_{n}/2\right)^{2d-2}$. Thus, $\nu_{ij}$ is the mean of at least $N_{n}$ independent subgaussian random variables of variance proxy $\sigma^{2}$, and by union bounding Lemma H.6 over all $n^{2}$ user-item pairs we have
$\mathbb{P}\left(\max_{(i,j):\bar{h}_{ij}\neq\star}\lvert\nu_{ij}\rvert\geq\sqrt{\frac{\sigma^{2}}{2cn^{2}\rho_{n}}\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\log\left(\frac{2n^{2}}{\delta}\right)}\right)\leq\delta.$ (92)
(2) Bounding $\Delta_{ij}$. Moreover, by Lemma A.1, all the estimated user-user and item-item distances have error bounded by $\varepsilon_{n}$ if
$\delta_{\operatorname{dist}}(n)=\frac{C}{\rho_{n}^{2}}\sqrt{\frac{d^{3}}{n}\log\left(\frac{e\,n}{\delta}\right)}\leq\varepsilon_{n}.$
@@ Appendix C Denoising
$\left\lvert f(\mathbf{u}_{k},\mathbf{v}_{l})-f(\mathbf{u}_{i},\mathbf{v}_{j})\right\rvert\leq 14LB\varepsilon_{n}.$ (93)
Combining the two bounds yields the desired bound. ∎
Appendix D Imputation
Pattern of observations and low-rank structure.
Once clustering &\&& denoising have been executed, the resulting matrix $\overline{\mathbf{H}}$ has still some missing entries, and its low-rank structure is a priori unclear. In Lemma D.1, we argue that $\overline{\mathbf{H}}$ has rank at least $d$, and describe the pattern its revealed entries. The theoretical tools we use towards proving Lemma D.1 exploit the connection between the geometry of the problem and the low-rank structure present in our model. These tools are presented in Appendix E.
Lemma D.1 (Patterns of observations).
Assume that $\theta(\cdot)$ and ϕ()italic-$\phi(\cdot)$ are non-degenerate with parameters $m,\kappa>0$. Let $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ be such that $\bar{h}_{i,j}=\star$, (i.e., $(i,j)$ is a non-revealed user-item cluster in $\overline{\mathbf{H}}$). Suppose we choose
$N_{n}=Cn^{2}\rho_{n}\varepsilon_{n}^{2d-2},$ (94)
where $C>0$ is an absolute constant sufficiently large. Then, there exists $\mathcal{I}\subset\mathcal{U}^{\star},\mathcal{J}\subseteq\mathcal{V}^{\star}$ with $|\mathcal{I}|=|\mathcal{J}|=d$, such that the event:
[…]
where $c,C,C^{\prime}>0$ are universal constants.
Proof of Lemma D.1.
Let $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ such that $\bar{h}_{ij}=\star$, and recall that $h_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})$. We start by invoking Proposition E.1, which ensures, following Algorithm 1, that there exists $\mathbf{x}_{1},\dots,\mathbf{x}_{d},$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ in $\mathcal{S}^{d-1}$, such that:
@@ Appendix D Imputation
Therefore, using Lemma B.1, there must exist $i_{1}^{\prime},\dots,i_{d}^{\prime}\in[n]$ such that $\|\mathbf{u}_{i_{k}^{\prime}}-\mathbf{z}_{k}\|\leq\varepsilon_{n}$ for all $k\in[d]$, which in turn ensures existence of $i_{1},\dots,i_{d}\in\mathcal{U}^{\star}$ such that $\|\mathbf{u}_{i_{k}}-\mathbf{z}_{k}\|\leq 7\varepsilon_{n}$. Finally, using property (i) together with triangular inequality ensures (97). Existence of $j_{1},\dots,j_{k}\in\mathcal{V}^{\star}$ satisfying (97) is proven similarly.
@@
$\|\mathbf{u}_{i}-\mathbf{v}_{j_{\ell}}\|_{2}\leq\sqrt{2}+8\varepsilon_{n},\quad\|\mathbf{u}_{i_{k}}-\mathbf{v}_{j_{\ell}}\|_{2}\leq\sqrt{2}+16\varepsilon_{n},\quad\|\mathbf{u}_{i_{k}}-\mathbf{v}_{j}\|_{2}\leq\sqrt{2}+8\varepsilon_{n}$ (99)
and by Lemma E.5 together with property $(iii)$, we obtain
$\sigma_{d}\left(\mathbf{H}_{\mathcal{I},\mathcal{J}}\right)=\sigma_{d}\left(\left(\theta(\mathbf{u}_{i_{k}})^{\top}\phi(\mathbf{v}_{j_{\ell}})\right)_{k,\ell\in[d]}\right)\geq\frac{m^{2}\kappa^{4}(1-\cos(\pi/12))\cos(\pi/12)}{4}$ (100)
@@
Since $\Sigma_{n}$ is an $\varepsilon_{n}$-net, using using (99), there exists $\mathbf{z}_{i},\mathbf{z}_{i_{1}},\dots,\mathbf{z}_{i_{d}}$ and $\mathbf{z}_{j},\mathbf{z}_{j_{1}},\dots,\mathbf{z}_{j_{d}}$ in $\Sigma_{n}$ such that for all $k,\ell\in[d]$, we have
$\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i}),\quad\mathbf{u}_{i_{k}}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i_{k}}),\quad\mathbf{v}_{j}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j}),\quad\mathbf{v}_{i_{\ell}}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i_{\ell}})$ (102)
@@
$\|\mathbf{z}_{i}-\mathbf{z}_{j_{\ell}}\|_{2}\leq\sqrt{2}+10\varepsilon_{n},\quad\|\mathbf{z}_{i_{k}}-\mathbf{z}_{j_{\ell}}\|_{2}\leq\sqrt{2}+18\varepsilon_{n},\quad\|\mathbf{z}_{i_{k}}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+10\varepsilon_{n},$ (103)
thus, given the choice of $N_{n}$, we may apply Lemma B.2 to conclude that the event:
$\forall i^{\prime}\in\mathcal{I},\forall j^{\prime}\in\mathcal{J},\qquad\bar{h}_{ij^{\prime}}\neq\star,\quad\bar{h}_{i^{\prime}j^{\prime}}\neq\star,\quad\bar{h}_{i^{\prime}j}\neq\star,$

### v3

@@ A.2.2 Observations
$\mathcal{E}_{1}$ =$=\left\{\sum_{k=1}^{n}\mathds{1}\{\mathbf{u}_{k}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i})\}\geq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right\},$
$\mathcal{E}_{2}$ =$=\left\{\sum_{\ell=1}^{n}\mathds{1}\{\mathbf{v}_{\ell}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j})\}\geq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right\},$
and note that according to Lemma 3 and union bounding over the two events we have
@@
$\left\lVert\mathbf{u}_{k}-\mathbf{v}_{l}\right\rVert_{2}$ $\leq\left\lVert\mathbf{z}_{i}-\mathbf{z}_{j}\right\rVert_{2}+\left\lVert\mathbf{u}_{k}-\mathbf{z}_{i}\right\rVert_{2}+\left\lVert\mathbf{v}_{l}-\mathbf{z}_{j}\right\rVert_{2}$
$\leq\sqrt{2}+(2+c)\varepsilon_{n}$
$\leq\sqrt{3}.$
@@
$\mathcal{E}_{3}=\left\{\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}^{(\mathbf{z}_{i},\mathbf{z}_{j})}\right\rvert\geq\frac{n^{2}\rho_{n}}{512}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right\},$
@@
$\mathbb{P}\left(\mathcal{E}_{3}^{c}|\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)$ ($\leq\left\lparen\frac{3}{\varepsilon_{n}}\right\rparen^{2d}\exp\left\lparen-\frac{\rho_{n}n^{2}}{2048}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right\rparen$ )
$\leq\delta,$
@@
$\rho_{n}n^{2}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\geq 2048\left(\log\left(\frac{1}{\delta}\right)+2d\log\left(\frac{3}{\varepsilon_{n}}\right)\right).$
@@
$\mathbb{P}\left(\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}^{(\mathbf{z}_{i},\mathbf{z}_{j})}\right\rvert<\frac{\rho_{n}n^{2}}{512}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right)$
$\qquad\qquad\qquad\leq\mathbb{P}\left(\mathcal{E}_{3}^{c}|\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)+\mathbb{P}(\mathcal{E}_{1}^{c}\cup\mathcal{E}_{2}^{c})$
$\qquad\qquad\qquad\leq 3\delta.$
A.3 Proof of Proposition 2
Now we are ready to prove Proposition 2. To start with, we can express the difference between $\bar{h}_{ij}$ and $h_{ij}$ for any $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ as $\bar{h}_{ij}-h_{ij}=\Delta_{ij}+\nu_{ij}$, where
$\Delta_{ij}$ $=\frac{1}{|\Omega_{ij}|}\sum_{(k,l)\in\Omega_{ij}}\left(f(\mathbf{u}_{k},\mathbf{v}_{l})-f(\mathbf{u}_{i},\mathbf{v}_{j})\right)$
$\nu_{ij}$ $=\frac{1}{|\Omega_{ij}|}\sum_{(k,l)\in\Omega_{ij}}\xi_{kl}$
(1) Bounding $\nu_{ij}$. Recall that an entry of $\overline{\mathbf{H}}$ is considered revealed if $\lvert\Omega_{ij}\rvert\geq N_{n}=c\,n^{2}\rho_{n}\left(\varepsilon_{n}/2\right)^{2d-2}$. Thus, $\nu_{ij}$ is the mean of at least $N_{n}$ independent subgaussian random variables of variance proxy $\sigma^{2}$, and by union bounding Lemma 15 over all $n^{2}$ user-item pairs we have that the event:
$\max_{(i,j):\bar{h}_{ij}\neq\star}\lvert\nu_{ij}\rvert\geq\sqrt{\frac{\sigma^{2}}{2cn^{2}\rho_{n}}\left(\frac{2}{\varepsilon_{n}}\right)^{2d-2}\log\left(\frac{2n^{2}}{\delta}\right)},$ (79)
holds with probability at most $\delta$
(2) Bounding $\Delta_{ij}$. Moreover, by Lemma 2, all the estimated user-user and item-item distances have error bounded by $\varepsilon_{n}$ if
$\delta_{\operatorname{dist}}(n)\triangleq\frac{C}{\rho_{n}^{2}}\sqrt{\frac{d^{3}}{n}\log\left(\frac{e\,n}{\delta}\right)}\leq\varepsilon_{n}.$
@@ A.3 Proof of Proposition 2
$\left\lvert f(\mathbf{u}_{k},\mathbf{v}_{l})-f(\mathbf{u}_{i},\mathbf{v}_{j})\right\rvert\leq 14LB\varepsilon_{n}.$
Combining the two bounds yields the desired bound.
Appendix B Proof of Proposition 3
Once clustering &\&& denoising have been executed, the resulting matrix $\overline{\mathbf{H}}$ has still some missing entries, and its low-rank structure is a priori unclear. In Proposition 3, we argue that $\overline{\mathbf{H}}$ has rank at least $d$, and describe the pattern of its revealed entries. The theoretical tools we use to prove Proposition 3 exploit the connection between the geometry of the problem and the low-rank structure present in our model. These tools are presented in Section B.2.
B.1 Proof of Proposition 3.
To start with, let us present an important result about the existence of unit vectors with some crucial properties that will prove useful for the proof of Proposition 3.
Proposition 5.
  • (i)
  • (ii)
  • (iii)
    and the following holds
We postpone the proof of Proposition 5 until the next section. With this we are now ready to present the proof of Proposition 5. Let $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ such that $\bar{h}_{ij}=\star$, and recall that $h_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})$. We start by invoking Proposition 5, which ensures, following Algorithm 1, that there exists $\mathbf{x}_{1},\dots,\mathbf{x}_{d},$ and $\mathbf{y}_{1},\dots,\mathbf{y}_{d}$ in $\mathcal{S}^{d-1}$, such that:
@@ Appendix B Proof of Proposition 3
Therefore, using Lemma 3, there must exist $i_{1}^{\prime},\dots,i_{d}^{\prime}\in[n]$ such that $\|\mathbf{u}_{i_{k}^{\prime}}-\mathbf{z}_{k}\|\leq\varepsilon_{n}$ for all $k\in[d]$, which in turn ensures existence of $i_{1},\dots,i_{d}\in\mathcal{U}^{\star}$ such that $\|\mathbf{u}_{i_{k}}-\mathbf{z}_{k}\|\leq 7\varepsilon_{n}$. Finally, using property (i) together with triangular inequality ensures (80). Existence of $j_{1},\dots,j_{k}\in\mathcal{V}^{\star}$ satisfying (80) is proven similarly.
@@
$\|\mathbf{u}_{i}-\mathbf{v}_{j_{\ell}}\|_{2}$ $\leq\sqrt{2}+8\varepsilon_{n},$
$\|\mathbf{u}_{i_{k}}-\mathbf{v}_{j_{\ell}}\|_{2}$ $\leq\sqrt{2}+16\varepsilon_{n},$
$\quad\|\mathbf{u}_{i_{k}}-\mathbf{v}_{j}\|_{2}$ $\leq\sqrt{2}+8\varepsilon_{n}$ (82)
and by Lemma 9 together with property $(iii)$, we obtain
$\sigma_{d}\left(\mathbf{H}_{\mathcal{I},\mathcal{J}}\right)$ $=\sigma_{d}\left(\left(\theta(\mathbf{u}_{i_{k}})^{\top}\phi(\mathbf{v}_{j_{\ell}})\right)_{k,\ell\in[d]}\right)$ (83)
$\geq\frac{m^{2}\kappa^{4}(1-\cos(\pi/12))\cos(\pi/12)}{4}$ (84)
@@
Since $\Sigma_{n}$ is an $\varepsilon_{n}$-net, using using (B.1), there exists $\mathbf{z}_{i},\mathbf{z}_{i_{1}},\dots,\mathbf{z}_{i_{d}}$ and $\mathbf{z}_{j},\mathbf{z}_{j_{1}},\dots,\mathbf{z}_{j_{d}}$ in $\Sigma_{n}$ such that for all $k,\ell\in[d]$, we have
$\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i}),\quad\mathbf{u}_{i_{k}}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i_{k}}),$
$\mathbf{v}_{j}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j}),\quad\mathbf{v}_{i_{\ell}}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i_{\ell}})$
@@
$\|\mathbf{z}_{i}-\mathbf{z}_{j_{\ell}}\|_{2}$ $\leq\sqrt{2}+10\varepsilon_{n},$
$\quad\|\mathbf{z}_{i_{k}}-\mathbf{z}_{j_{\ell}}\|_{2}$ $\leq\sqrt{2}+18\varepsilon_{n},$
$\quad\|\mathbf{z}_{i_{k}}-\mathbf{z}_{j}\|_{2}$ $\leq\sqrt{2}+10\varepsilon_{n},$
thus, given the choice of $N_{n}$, we may apply Lemma 4 to conclude that the event:
$\forall i^{\prime}\in\mathcal{I},\forall j^{\prime}\in\mathcal{J},\ \ \bar{h}_{ij^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime}j^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime}j}\neq\star,$